Astrid1:

Check on the 1st one? ssbelow Idr any of these so help me. I've got plenty more, i need help w/ btw. xd

Astrid1:

XLynx:

What grade are you in?

Astrid1:

8th?

XLynx:

Interesting, did you take the test to get into Algebra 1?

Astrid1:

No?

Astrid1:

Can we just stay on topic? Can you help me?

mykenzie:

what one do u need help on

Astrid1:

All of them-

XLynx:

Okay then. Checking her answers @mykenzie. Sorry for getting distracted, I just wanted a bit of backstory.

mykenzie:

ok

Astrid1:

\( (a^b)^c = a^{bc}\) \( \dfrac{x^y}{x^z} = x^{y-z}\) \( a^b \times a^c = a^{b+c}\) Is it just these steps for them?

snowflake0531:

Yes And btw, the first one is wrong._.

Astrid1:

I thought sum was wrong w. #1. What did I do wrong?

mykenzie:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 Yes And btw, the first one is wrong._. \(\color{#0cbb34}{\text{End of Quote}}\) yes it is

mykenzie:

i think that she did a wrong

snowflake0531:

To the contrary, I think that she did A right But, D is not right

snowflake0531:

Let's first look at B It's 10 to the power of -3, so just cross out 3 zeroes, does it equal 6 with 5 zeroes?

Astrid1:

Yes?

snowflake0531:

Okay And for C, is it true where 6 is the head with 5 more zeroes?

Astrid1:

Yes?

snowflake0531:

._. lel, are you just guessing?

Astrid1:

(._.') Maybe?

snowflake0531:

LOL anyways, so we ended up with A, B, and C correct xd

Astrid1:

xd Okayy...

snowflake0531:

And now look at problem 2, and remember the exponent properties you earlier copied over

Astrid1:

I know for number two one of the ones are b, correct?

snowflake0531:

no.... reread the exponent properties-

Astrid1:

e.e Okayyyyy.....

Astrid1:

I can see my mistake.I added the exponents instead of multiplying.

snowflake0531:

Yep And so now, relook at all the choices which ones do you think are correct?

Astrid1:

Wouldn't it just be D?

snowflake0531:

Why would D be correct...., D would make it to the power of -8... but we want 8

Astrid1:

OOoooooooooffffff, This is so hard. ;c

snowflake0531:

._. Just relook at those properties LEt's first look at A what would the exponent be?

Astrid1:

Thanks for ya partial help, I ended up finishing it on my own.

Quicksilver0125:

im in 8th grade 2